OpenStudy (eric_d):

(x^2+5^2)^2 = x^2+5^2+2(x^2)(5) =11x^2+25 Is this working correct @ikram002p

mathslover (mathslover):

Eric, recheck it. \((a+b)^2= a^2 + b^2 +2ab\) here, you have a = x^2 and b = 5^2

OpenStudy (ikram002p):

wait u know the rule , just set up it correctly :D

OpenStudy (anonymous):

in the first and second terms, x and 5 are raised to the fourth power not the second. and in the last term five should be squared. so you will end up with x^4+5^4+2(x^2)(5^2)

OpenStudy (eric_d):

Understood

OpenStudy (eric_d):

Thnx

OpenStudy (ikram002p):

but is ur Qn asking for expanding or finding the inverse?

OpenStudy (eric_d):

It's nt a complete question I managed to solve it nw :)

OpenStudy (ikram002p):

ohk :D good luck !